OpenStudy (chunkymonkay):

which answer shows this question in standard form? 6-2(x-y)=-3x+5

OpenStudy (chunkymonkay):

A. x+2y=-1 B.x+2y=1 C.x-2y=1 D.-5x+2y=-1

OpenStudy (chunkymonkay):

@dtan5457 I think its d but im not sure

OpenStudy (dtan5457):

Did you distribute the parenthesis?

OpenStudy (chunkymonkay):

yeah when I distributed them I got 6-2x+2y=3x+5

OpenStudy (dtan5457):

This is weird, I;m not getting any of your answer choices, can you check again?

OpenStudy (chunkymonkay):

yeah I wrote them down all correctly

OpenStudy (dtan5457):

and the question?

OpenStudy (chunkymonkay):

yep that's correct too

OpenStudy (dtan5457):

6-2(x-y)=-3x+5 6x-6y-2x+2y=-3x+5 4x-4y=-3x+5 7x-4y=5<<<not an answer choice

OpenStudy (chunkymonkay):

I did it like this. 6-2(x-y)=-3x+5 6-2x+2y=-3x+5 5x-2y=-1

OpenStudy (dtan5457):

lol. you just solved your own problem. I looked at this like (6-2)(x-y), in which you will foil. 6-2(x-y) you factored it correctly. you did it right . lol my bad

OpenStudy (chunkymonkay):

really!! yay

OpenStudy (dtan5457):

yeah. gotta remember how important parenthesis are...sigh

OpenStudy (chunkymonkay):

lol yeah guess im not that supid after all

OpenStudy (dtan5457):

every one makes mistakes but you were never stupid in the first place

OpenStudy (dtan5457):

anymore questions? ill try to make dumb mistakes like that

OpenStudy (dtan5457):

*not

OpenStudy (dtan5457):

lol

OpenStudy (chunkymonkay):

haha ill tag you in some more questions that I have

OpenStudy (dtan5457):

ok, i'll be on for the next 40 minutes.